# Weinstein exactness of nearby Lagrangians and the Lagrangian $C^{0}$ flux conjecture

### v2

@@ 1.1.2 Lagrangian flux conjectures
We now move on to another conjecture about Lagrangian submanifolds.
Conjecture C (Lagrangian $C^{0}$ flux conjecture)
@@
As far as the authors know, this version of the conjecture has not been studied previously — we will talk about its $C^{1}$ cousin, which has been studied, below. The name that we give it here is in analogy to the famous $C^{0}$ flux conjecture for Hamiltonian diffeomorphisms, which states that the group $\mathrm{Ham}(M)$ of Hamiltonian diffeomorphisms of a closed manifold $M$ is $C^{0}$ closed in the identity component $\mathrm{Symp}_{0}(M)$ of the group of symplectomorphisms of $M$. This conjecture is only known to hold in some fairly specific case [LMP98, Buh15]. This is in stark contrast with its $C^{1}$ cousin, which is known to hold in full generality [Ono06].
We note that, similarly to Conjecture A above, the Lagrangian $C^{0}$ flux conjecture does not imply the one for Hamiltonian diffeomorphisms. Indeed, suppose that $\{\varphi_{i}\}\subseteq\mathrm{Ham}(M)$ $C^{0}$-converges to $\psi\in\mathrm{Symp}_{0}(M)$. Then, all that the Lagrangian flux conjecture ensures is that there is some Hamiltonian diffeomorphism $\Phi$ of $M\times M$ such that $\mathrm{graph}(\psi)=\Phi(\Delta)$, where $\Delta\subseteq M\times M$ is the diagonal. However, we cannot be sure that $\Phi$ can be chosen of the form $\mathds{1}\times\varphi$ for some $\varphi\in\mathrm{Ham}(M)$, which is what the flux conjecture for Hamiltonian diffeomorphisms would require.
To study this conjecture, we can use the techniques developed to prove Theorems 3 and 4. In fact, they imply the following continuity result.
@@
Moreover, if $L_{i}$ is $H$-exact in a Weinstein neighbourhood $\mathcal{W}(L)$ for $i$ large, then $\tau_{i}\equiv\tau$ for larger $i$, and $\tau$ is the $H$-rationality constant of $L$. This is in particular the case if the $L_{i}$’s respect the hypotheses of Theorem 3.
By tame, we mean that $M$ admits an almost structure $J$ making $g_{J}:=\omega(\cdot,J\cdot)$ into a complete Riemannian metric whose sectional curvature is bounded and whose injectivity radius is bounded away from zero.
@@
Before moving on to corollaries of this result, note that, in the formulation above, one could also ask for $C^{0}$-closure of $\mathrm{Ham}(M)$ in larger groups than $\mathrm{Symp}_{0}(M)$, most notably in $\mathrm{Symp}(M)$ or $\mathrm{Diff}(M)$. Here, $\mathrm{Symp}(M)$ denotes the group of symplectomorphisms of $M$ and $\mathrm{Diff}(M)$, the group of its diffeomorphisms. For $\mathrm{Symp}(M)$, this is harder to prove, since it is not known if $\mathrm{Symp}_{0}(M)$ is $C^{0}$-closed in $\mathrm{Symp}(M)$. However, the celebrated result from Gromov [Gro85] and Eliashberg [Eli87] shows that $\mathrm{Symp}(M)$ is $C^{0}$-closed in $\mathrm{Diff}(M)$. Therefore, the closure in $\mathrm{Symp}(M)$ is equivalent to that in $\mathrm{Diff}(M)$.
Following this logic, we can replace $\mathcal{L}\mathrm{Lag}(L)$ in the above with larger spaces. Most notably, we will also be interested in the spaces $\mathsf{SMan}(L)$, of all submanifolds of $M$ with the same diffeomorphism type as $L$, and $\mathsf{SMan}_{n}$, of all $n$-dimensional submanifolds of $M^{2n}$. By Theorem 6, closure of $\mathcal{L}{\mathrm{Ham}}(L)$ in the two latter spaces is equivalent to closure in the subspace formed by Lagrangian submanifolds.
@@
Let $L$ be a $H$-rational Lagrangian submanifold of $M$ with $H$-rationality constant $\tau$. There is some $r_{0}>0$ with the following property. Assume that $L^{\prime}\in\mathcal{L}(\tau)$ is a Lagrangian included in a Weinstein neighbourhood $\mathcal{W}_{r}(L)$ of size $r\in(0,r_{0}]$ such that $L^{\prime}$ is $H$-exact in $\mathcal{W}_{r}(L)$. Then, there is a symplectic isotopy $\{\psi_{t}\}_{t\in[0,1]}$ of $M$ such that $\psi_{1}(L^{\prime})$ is exact in $\mathcal{W}_{r}(L)$. Furthermore, the size of the isotopy is controlled by $r$.
@@
Let $L$ be a $H$-rational Lagrangian in a tame symplectic manifold $M$. Suppose that
  1. (i)
    $L$ satisfies the hypotheses of Theorem 3 or $H_{1}(L;\mathds{R})=0$
  2. (ii)
    and the nearby Lagrangian conjecture holds in $T^{*}L$.
Then, $\mathcal{L}{\mathrm{Ham}}(L)$ and $\mathcal{L}{\mathrm{Symp}_{0}}(L)$ are Hausdorff-closed in $\mathsf{SMan}(L)$.
Remarks 3.
Note that the tameness condition on $M$ can be dropped if one is only interested in closedness in the subspace $\mathcal{L}(L)$ of $\mathsf{SMan}(L)$ consisting of the Lagrangians. Indeed, tameness is only used to ensure that the limit of Lagrangians is still a Lagrangian.
Likewise, one can, in some contexts, upgrade from $\mathsf{SMan}(L)$ to $\mathsf{SMan}_{n}$. For example, if $n=2$, this is the case. Indeed, any $H$-exact Lagrangian in the cotangent bundle of a surface has the same diffeomorphism type as that surface (see Lemma 24 below). This is a nontrivial update: Polterovich [Pol93] constructed Lagrangian tori in the cotangent bundle of any flat manifold; these tori can be made to be arbitrarily close to the zero-section. We discuss these examples in more details at the very end of Section 3.2.
@@
A natural variant of Conjecture C is obtained by replacing closedness in the Hausdorff metric with closedness in the $C^{1}$ topology. We call this the Lagrangian $C^{1}$ flux conjecture.
@@
The reason that we don’t need hypothesis (i) of Corollary 8 is because Lagrangians with are $C^{1}$-close of $L$ are graphs of 1-forms in $\mathcal{W}(L)$, and graphs are necessarily $H$-exact. Likewise, hypothesis (ii) is not needed since exact graphs are Hamiltonian isotopic to the zero-section in $\mathcal{W}(L)$. Note that $C^{1}$-close Lagrangians are necessarily diffeomorphic so that closure in $\mathsf{SMan}_{n}$ is the same as closure in $\mathsf{SMan}(L)$.
@@ 1.2 Examples
We give a few examples where Conjecture B follows from the results above. Note that, as long as $\omega(H_{2}(M;\mathds{Z}))$, an arbitrarily $C^{1}$-small perturbation of a Lagrangian will be $H$-rational. Therefore, we do not care if the precise examples below are $H$-rational or not.
We start by giving a few examples of Lagrangians satisfying Case (a) of Theorem 3 and natural situations when they are displaceable.
First of all, note that for any Liouville domain $V$, $W=V\times\mathds{D}$ is such that $SH(W)=0$, since that domain is displaceable in its completion. In particular, Case (a) covers the case $L=Q\times S^{1}$ with $H_{1}(Q;\mathds{R})=0$. Note that these are displaceable since $\mathds{D}$ is displaceable in $\mathds{C}$.
If $TQ\otimes\mathds{C}$ is additionally assumed to be trivial, then $L$ embeds as a Lagrangian in $\mathds{C}^{n}$ by the Gromov–Lees $h$-principle [Gro70, Lee76] and a result of Audin, Lalonde, and Polterovich [ALP94]. In particular, $L=S^{n-1}\times S^{1}$, $n\geq 3$ is an example of such a manifold. In particular, we have examples of $L$ satisfying Case (a) in any symplectic manifold of dimension at least 6. In fact, any symplectic manifold containing a displaceable Lagrangian $S^{n}$, e.g. the full flag manifold of $\mathds{C}^{3}$ [Pab15], will contain such Lagrangians, because $T^{*}S^{n}$ has Lagrangians $S^{n-1}\times S^{1}$ arbitrarily close to the zero-section.
In another direction, Ekholm, Eliashberg, Murphy, and Smith [EEMS13] showed that, given any 3-manifold $Q$, $L=Q\#(S^{1}\times S^{2})$ embeds as a Lagrangian in $\mathds{C}^{3}$. But, by the van Kampen theorem, $\pi_{1}(L)=\pi_{1}(Q)*\pi_{1}(S^{1}\times S^{2})$, so that $H_{1}(L;\mathds{R})=H_{1}(Q;\mathds{R})\oplus\mathds{R}$. Therefore, Case (i) covers $L=Q\#(S^{1}\times S^{2})$ with $H_{1}(Q;\mathds{R})=0$, e.g. $Q$ can be a (connected sum of) lens spaces.
To resume the above discussion, we have the following.
@@
Conjecture B holds for displaceable $H$-rational Lagrangians of the form $Q\times S^{1}$ or $Q\#(S^{1}\times S^{2})$ with $H_{1}(Q;\mathds{R})=0$ and, in the latter case, $\dim Q=3$.
There are of course examples which do not fit within this pattern. For example, it is well known that the Lagrangian Grassmannian $\Lambda_{n}$ admits a Lagrangian embedding in $\mathrm{Sym}(\mathds{C}^{n})=\mathds{C}^{n(n+1)/2}$ (see, for example, [ALP94]). This is an example of Case (a), since $\pi_{1}(\Lambda_{n})=\mathds{Z}$.
When it comes to Case (b) of Theorem 3, there is one main example: the Lagrangian Klein bottle in $S^{2}\times\mathds{C}$. It is obtained from the usual Lagrangian Klein bottle in $S^{2}\times S^{2}$ (see, for example, [Eva22]) by removing a point on the second copy of $S^{2}$ and identifying $S^{2}\setminus\{pt\}$ with $\mathds{D}\subseteq\mathds{C}$. Again, it is displaceable, because $\mathds{D}$ is. In fact, the Klein bottle can even be made to be monotone. To resume, we have the following.
@@
Conjecture B holds for displaceable $H$-rational Lagrangian Klein bottles. There exist such Lagrangians — and a monotone one — in $S^{2}\times\mathds{C}$.
To conclude with Conjecture B, we go back to Remark 2: using different methods, we can prove that the conjecture holds for rational 2-tori in simply connected Darboux charts. Using Theorem C of [RGI16], we, in particular, get the following result.
Corollary 12
Conjecture B holds for displaceable rational 2-tori in $\mathds{C}^{2}$, $S^{2}\times S^{2}$, $\mathds{C}P^{2}$, and blow-ups of $\mathds{C}P^{2}$.
We conclude with one additional case when we can establish Conjecture A: when $L$ is a 2-sphere or a projective plane. Indeed, any other such Lagrangian in $\mathcal{W}(L)$ is then automatically exact in that neighbourhood, so there is no need for Theorems 3 or 4. Since the NLC is known to hold in $T^{*}S^{2}$ [Hin04] and $T^{*}\mathds{R}P^{2}$ [HPW16], we thus directly get the following.
Corollary 13
Conjectures A and C hold for Lagrangian 2-spheres or projective planes.
@@ 1.3 Applications
We end this introduction with several applications of our results. These are divided in four parts: additional rigidity results on Lagrangians with regards to Hamiltonian diffeomorphisms, further study on the local topology of $\mathcal{L}{\mathrm{Ham}}(L)$, new results on the space of (H-)rational Lagrangians with a fixed rationality constant, and some computations of numerical invariants. The next to last part has further implications when it comes to the space of all Lagrangians of a given symplectic manifold.
This last part of the introduction is intended to be almost completely self-contained, using the results above as black boxes (except for a couple of references to further results when finer technical variants are needed.)
@@
There is a natural variant of Conjecture B where we ask not that $L^{\prime}$ be close to $L$, but rather that the Hamiltonian diffeomorphism sending $L$ to $L^{\prime}$ be small. More precisely, we can make the following conjecture.
Conjecture D
@@
However, this is not expected to be the case for the spectral metric, that is, $C^{0}$-small Hamiltonian diffeomorphisms should also have small spectral norm. More precisely, Conjecture D follows from the fact that Lagrangians have positive spectral displacement energy [AAC23] in the cases where it is known that the spectral metric is $C^{0}$-continuous, i.e. when $M$ is $\mathds{C}^{n}$ [Vit92], a closed surface [Sey13], closed and symplectically aspherical [BHS21], $\mathds{C}P^{n}$ [She22], or closed and negative monotone [Kaw22].
In the context of this paper, Conjecture D is implied by Conjecture A or by the Hamiltonian version of Conjecture B above when it holds. However, it turns out to be much easier to prove than either one of these conjectures. More precisely, we have the following lemma.
@@
Take a Riemannian metric $g$ on $M$ which corresponds to a Sasaki metric on $T^{*}L$ on a Weinstein neighbourhood $\mathcal{W}(L)$. With such a metric, the geodesics starting at $L$ and going to $L^{\prime}=\psi(L)$ stay in $\mathcal{W}(L)$ (see Lemma A.4 of [Cha24] for example). Therefore, if we assume that $\delta$ is smaller than the injectivity radius $r_{\mathrm{inj}}(TM|_{L})$ of the Riemannian exponential of $g$ restricted to $TM|_{L}$, we get for every $x\in L$ a unique geodesic $\gamma_{x}:[0,1]\to M$ such that $\gamma_{x}(0)=x$, $\gamma_{x}(1)=\psi(x)$, and $\gamma_{x}([0,1])\subseteq\mathcal{W}(L)$. Moreover, $\gamma_{x}$ smoothly depends on $x$. Therefore, $(x,t)\mapsto\gamma_{x}(t)$ defines a smooth homotopy in $\mathcal{W}(L)$ from the inclusion $\iota:L\hookrightarrow\mathcal{W}(L)$ to $\varphi\iota$. Since $\iota$ is a homotopy equivalence, then so must be $\varphi\iota$. In particular, $H_{2}(\mathcal{W}(L),L^{\prime})=0$, and $L^{\prime}$ is $H$-exact.
@@
Conjecture D holds for $H$-rational Lagrangians.
Likewise, we get a rigidity result for sequences of Hamiltonian or symplectic diffeomorphisms from Theorem 6 and Corollary 8.
@@
Local contractibility of $\mathcal{L}{\mathrm{Ham}}(L)$
Even though we need the NLC in all cases where we can prove Conjecture A, its implication that $\mathcal{L}{\mathrm{Ham}}(L)$ is locally path connected turns out to be easier to prove. More precisely, we get the following.
@@
Suppose that $L$ is $H$-rational and respects the hypotheses of Theorem 3. Then $\mathcal{L}{\mathrm{Ham}}(L)$ is locally contractible in the Hausdorff metric.
Note that this is not quite the type of results we mentioned earlier in the introduction. Indeed, we do not claim that the Hausdorff-continuous path from $L^{\prime}$ to $L$ in a small neighbourhood of $L$ is generated by an actual Hamiltonian isotopy, simply that it stays at all time in $\mathcal{L}{\mathrm{Ham}}(L)$.
@@
Note that it suffices to prove this statement at $L$. Fix a Weinstein neighbourhood $\Psi:D^{*}_{r}L\to\mathcal{W}(L)$ as given by the conclusions of Theorems 3 and 4. Then, every $L^{\prime}\in\mathcal{L}{\mathrm{Ham}}(L)$ which is in $\mathcal{W}(L)$ is exact in that neighbourhood. We can thus take
@@
to be the contraction. Indeed, exactness in $\mathcal{W}(L)$ ensures that this is a Hamiltonian isotopy for all $t>0$. But exactness also implies that the projection $\Psi^{-1}(L^{\prime})\to T^{*}L\to L$ is a homotopy equivalence [AK18]. In particular, that projection must be surjective, otherwise $H_{n}(L^{\prime})\to H_{n}(L)\neq 0$ would be zero. Therefore, $L^{\prime}$ being close to $L$ implies that $L$ is close to $L^{\prime}$. This mean that the Hausdorff limit of $\Psi(t\Psi^{-1}(L^{\prime}))$ as $t\to 0$ is precisely $L$, i.e. the above contraction is indeed continuous.
@@
From Theorems 3 and 4, we get the following.
@@
Let $L$ be a $H$-rational Lagrangian in a tame symplectic manifold, and denote by $\tau$ its $H$-rationality constant. Then, $\mathcal{L}{\mathrm{Symp}_{0}}(L)$ is open in $\mathcal{L}(L,\tau)$ in the $C^{1}$ topology. If moreover $L$ respects the hypotheses of Theorem 3 or $H_{1}(L;\mathds{R})=0$ and the NLC holds on $T^{*}L$, then the same holds in the Hausdorff topology.

### v3

@@ 1.1.3 Lagrangian flux conjectures
We now move on to another speculation about Lagrangian submanifolds.
Speculation D (Lagrangian $C^{0}$ flux conjecture)
@@
As far as the authors know, this version of the conjecture has not been studied previously — we will talk about its $C^{1}$ cousin, which has been studied, below. The name that we give it here is in analogy to the famous $C^{0}$ flux conjecture for Hamiltonian diffeomorphisms, which states that the group $\mathrm{Ham}(M)$ of Hamiltonian diffeomorphisms of a closed symplectic manifold $M$ is $C^{0}$ closed in the identity component $\mathrm{Symp}_{0}(M)$ of the group of symplectomorphisms of $M$. This conjecture is only known to hold in some fairly specific cases [LMP98, Buh15] and further results in this direction will appear in [AS24], which still do not resolve this question completely. This is in stark contrast with its $C^{1}$ cousin, which is known to hold in full generality [Ono06].
We note that, similarly to Speculation A above, the Lagrangian $C^{0}$ flux conjecture does not imply the one for Hamiltonian diffeomorphisms. Indeed, suppose that $\{\varphi_{i}\}\subseteq\mathrm{Ham}(M)$ $C^{0}$-converges to $\psi\in\mathrm{Symp}_{0}(M)$. Then, all that the Lagrangian flux conjecture ensures is that there is some Hamiltonian diffeomorphism $\Phi$ of $M\times M$ such that $\mathrm{graph}(\psi)=\Phi(\Delta)$, where $\Delta\subseteq M\times M$ is the diagonal. However, we cannot be sure that $\Phi$ can be chosen of the form $\mathds{1}\times\varphi$ for some $\varphi\in\mathrm{Ham}(M)$, which is what the flux conjecture for Hamiltonian diffeomorphisms would require.
To study this speculation, we can use the techniques developed to prove Theorems 4 and 5. In fact, they imply the following continuity result.
@@
Moreover, if $L_{i}$ is $H$-exact in a Weinstein neighbourhood $\mathcal{W}(L)$ for $i$ large, then $\lim\tau_{i}$ exists and is the $H$-rationality constant of $L$. This is in particular the case if the $L_{i}$’s respect the hypotheses of Theorem 4.
By tame, we mean that $M$ admits an almost complex structure $J$ making $g_{J}:=\omega(\cdot,J\cdot)$ into a complete Riemannian metric whose sectional curvature is bounded and whose injectivity radius is bounded away from zero.
@@
Before moving on to corollaries of this result, note that, in the formulation above, one could also ask for $C^{0}$-closure of $\mathrm{Ham}(M)$ in larger groups than $\mathrm{Symp}_{0}(M)$, most notably in $\mathrm{Symp}(M)$.
Following this logic, we can replace $\mathcal{L}\mathrm{Lag}(L)$ in Speculation D with larger spaces. Most notably, we will also be interested in the spaces $\mathsf{SMan}(L)$, of all submanifolds of $M$ with the same diffeomorphism type as $L$, and $\mathsf{SMan}_{n}$, of all $n$-dimensional submanifolds of $M^{2n}$. By Theorem 7, closure of $\mathcal{L}{\mathrm{Ham}}(L)$ in the two latter spaces is equivalent to closure in the subspace formed by Lagrangian submanifolds.
@@
Let $L$ be a $H$-rational Lagrangian submanifold of $M$ with $H$-rationality constant $\tau$. There exists $r_{0}>0$ with the following property. Assume that $L^{\prime}\in\mathcal{L}(\tau)$ is a Lagrangian included in a Weinstein neighbourhood $\mathcal{W}_{r}(L)$ of radius $r\in(0,r_{0}]$ such that $L^{\prime}$ is $H$-exact in $\mathcal{W}_{r}(L)$. Then, there is a symplectic isotopy $\{\psi_{t}\}_{t\in[0,1]}$ of $M$ such that $\psi_{1}(L^{\prime})$ is exact in $\mathcal{W}_{r}(L)$. Furthermore, the size of the isotopy is controlled by $r$.
@@
Suppose that $L$ satisfies the hypotheses of Theorem 2 or that $H_{1}(L;\mathds{R})=0$. If the nearby Lagrangian conjecture holds in $T^{*}L$, then we have the following. Suppose that $L^{\prime}$ is a $H$-rational Lagrangian diffeomorphic to $L$. Its Hamiltonian orbit $\mathcal{L}{\mathrm{Ham}}(L^{\prime})$ and symplectic orbit $\mathcal{L}{\mathrm{Symp}_{0}}(L^{\prime})$ are Hausdorff-closed in $\mathcal{L}(L)$.
In fact, we can prove the equivalence of the NLC and the Lagrangian $C^{0}$ flux conjecture in some cases — we refer the interested reader to Section 5.3.
Remark 4.
Note that one can upgrade from $\mathcal{L}(L)$ to $\mathsf{SMan}(L)$ under the additional condition that the ambient symplectic manifold is tame.
Furthermore, one can upgrade from $\mathsf{SMan}(L)$ to $\mathsf{SMan}_{n}$ in some contexts as, for example, if $n=2$. Indeed, any $H$-exact Lagrangian in the cotangent bundle of a surface has the same diffeomorphism type as that surface (see Lemma 24 below). This is a nontrivial update: Polterovich [Pol93] constructed Lagrangian tori in the cotangent bundle of any flat manifold; these tori can be made to be arbitrarily close to the zero-section. We discuss these examples in more details at the very end of Section 3.2.
@@
A natural variant of Speculation D is obtained by replacing closedness in the Hausdorff topology with closedness in the $C^{1}$ topology. We call this the Lagrangian $C^{1}$ flux conjecture.
@@
The reason there is no restriction on the diffeomorphism type as in Corollary 9 is because Lagrangians which are $C^{1}$-close to $L$ are graphs of 1-forms in $\mathcal{W}(L)$, and graphs are necessarily $H$-exact. Likewise, the NLC is not needed since exact graphs are Hamiltonian isotopic to the zero-section in $\mathcal{W}(L)$. Note that $C^{1}$-close Lagrangians are necessarily diffeomorphic so that closure in $\mathsf{SMan}_{n}$ is the same as closure in $\mathsf{SMan}(L)$.
@@ 1.2 Examples
We give a few examples where Speculation B follows from the results above. We will only specify a Lagrangian submanifold in a symplectic manifold without worrying about it being rational or not, since given an irrational Lagrangian submanifold $L^{\prime}$ of a rational symplectic manifold, we can construct a nearby rational Lagrangian $L$ by an arbitrarily $C^{1}$-small perturbation.
First, from Theorem 2 we know that Conjecture C holds for closed connected manifolds $L$ which are covered by $L_{0}\times\mathds{T}^{m}$, where $H_{1}(L_{0};\mathds{R})=0$ and $m$ is allowed to be equal to zero. For example, this includes every manifold admitting a flat metric by Bieberbach’s theorem. Curiously, such manifolds also include every mapping torus $L$ of a diffeomorphism $f\in\operatorname{\mathrm{Diff}}(L_{0})$ of a simply connected manifold $L_{0}$ that has finite order in the smooth mapping class group, i.e. $f^{k}$ is smoothly isotopic to $\operatorname{\mathrm{id}}$ for some integer $k.$ Indeed, the multiplication by $k$ map $S^{1}\to S^{1}$ pulls the bundle $L\to S^{1}$ to a bundle diffeomorphic to $L_{0}\times S^{1}$. Furthermore, note that the class of manifolds covered by $L_{0}\times\mathds{T}^{m}$ (where $m$ and $L_{0}$ are not fixed) is also closed under products, so that Conjecture C also holds for all products of the above examples.
Second, we give a few examples of Lagrangian submanifolds $L$ where Speculation B holds. By Corollary 6, this includes those $L$ such that $H_{1}(L;\mathds{R})=\mathds{R}$ which admit embeddings to a Liouville domains $W$ with $SH(W)=0.$ For instance if $TL_{0}\otimes\mathds{C}$ is trivial and $H_{1}(L_{0};\mathds{R})=0$, then $L=L_{0}\times S^{1}$ embeds as a Lagrangian in $\mathds{C}^{n}$ by the Gromov–Lees $h$-principle [Gro70, Lee76] and a result of Audin, Lalonde, and Polterovich [ALP94]. In another direction, Ekholm, Eliashberg, Murphy, and Smith [EEMS13] showed that, given any 3-manifold $L_{0}$, $L=L_{0}\#(S^{1}\times S^{2})$ embeds as a Lagrangian in $\mathds{C}^{3}$. But, by the van Kampen theorem, $\pi_{1}(L)=\pi_{1}(L_{0})*\pi_{1}(S^{1}\times S^{2})$, so that $H_{1}(L;\mathds{R})=H_{1}(L_{0};\mathds{R})\oplus\mathds{R}$. Therefore, Corollary 6 covers $L=L_{0}\#(S^{1}\times S^{2})$ with $H_{1}(L_{0};\mathds{R})=0$, e.g. $L_{0}$ can be a (connected sum of) lens spaces. Finally, the Lagrangian Grassmannian $L=\Lambda_{n}$ admits a Lagrangian embedding in $\mathrm{Sym}(\mathds{C}^{n})=\mathds{C}^{n(n+1)/2}$ (see, for example, [ALP94]), and $\pi_{1}(\Lambda_{n})=\mathds{Z}$.
Specializing to the case of the 2-torus, where the NLC is known to hold [RGI16], we get the following.
@@
Speculations A and D hold for all $H$-rational Lagrangian 2-tori.
Again, recall that the class for which Corollary 6 holds is closed under covering. But note that the Klein bottle admits a displaceable Lagrangian embedding in $S^{2}\times\mathds{C}$. It is obtained from the usual Lagrangian Klein bottle in $S^{2}\times S^{2}$ (see, for example, [Eva22] and [AE24]) by removing a point on the second copy of $S^{2}$ and identifying $S^{2}\setminus\{pt\}$ with $\mathds{D}\subseteq\mathds{C}$. Again, it is displaceable, because $\mathds{D}$ is. In fact, the Klein bottle can even be made to be monotone.
We conclude with one additional case when we can establish Speculation A: when $L$ is a 2-sphere or a projective plane. Indeed, any other such Lagrangian in $\mathcal{W}(L)$ is then automatically exact in that neighbourhood, so there is no need for Theorems 4 or 5. Since the NLC is known to hold in $T^{*}S^{2}$ [Hin04] and $T^{*}\mathds{R}P^{2}$ [HPW16], we thus directly get the following.
@@
Speculations A and D hold for Lagrangian 2-spheres or projective planes.
@@ 1.3 Applications
We end this introduction with applications of our results. These are divided in four parts: additional rigidity results on Lagrangians with regards to Hamiltonian diffeomorphisms, further study on the local topology of $\mathcal{L}{\mathrm{Ham}}(L)$, new results on the space of (H-)rational Lagrangians with a fixed rationality constant, and some computations of numerical invariants. The next to last part has further implications when it comes to the space of all Lagrangians of a given symplectic manifold.
Except for a couple of references to further results, this last part of the introduction is intended to be self-contained; we directly give the proofs when these are not direct.
@@
There is a natural variant of Speculation B where we ask not that $L^{\prime}$ be close to $L$, but rather that the Hamiltonian diffeomorphism sending $L$ to $L^{\prime}$ be small. More precisely, we can make the following conjecture.
Conjecture E
@@
However, this is not expected to be the case for the spectral metric, that is, $C^{0}$-small Hamiltonian diffeomorphisms should also have small spectral norm. More precisely, Conjecture E follows from the fact that Lagrangians have positive spectral displacement energy [AAC23] in the cases where it is known that the spectral metric is $C^{0}$-continuous, i.e. when $M$ is $\mathds{C}^{n}$ [Vit92], a closed surface [Sey13], closed and symplectically aspherical [BHS21], $\mathds{C}P^{n}$ [She22], or closed and negative monotone [Kaw22] or symplectically Calabi-Yau [SW].
In the context of this paper, Conjecture E is implied by Speculation A or by the Hamiltonian version of Speculation B above when it holds. However, it turns out to be much easier to prove than either one of these speculations. This is because we have the following lemma.
@@
Take a Riemannian metric $g$ on $M$ which corresponds to a Sasaki metric on $T^{*}L$ on a Weinstein neighbourhood $\mathcal{W}(L)$. With such a metric, the geodesics starting at $L$ and going to $L^{\prime}=\psi(L)$ stay in $\mathcal{W}(L)$ (see Lemma A.4 of [Cha24] for example). Therefore, if we assume that $\delta$ is smaller than the injectivity radius $r_{\mathrm{inj}}(TM|_{L})$ of the Riemannian exponential of $g$ restricted to $TM|_{L}$, we get for every $x\in L$ a unique geodesic $\gamma_{x}:[0,1]\to M$ such that $\gamma_{x}(0)=x$, $\gamma_{x}(1)=\psi(x)$, and $\gamma_{x}([0,1])\subseteq\mathcal{W}(L)$. Moreover, $\gamma_{x}$ smoothly depends on $x$. Therefore, $(x,t)\mapsto\gamma_{x}(t)$ defines a smooth homotopy in $\mathcal{W}(L)$ from the inclusion $\iota:L\hookrightarrow\mathcal{W}(L)$ to $\psi\iota$. Since $\iota$ is a homotopy equivalence, then so must be $\psi\iota$. In particular, $H_{2}(\mathcal{W}(L),L^{\prime})=0$, and $L^{\prime}$ is $H$-exact.
@@
Conjecture E holds for all $H$-rational Lagrangians.
In fact, at the price of making the constant $\delta$ depend on the size of the compact support of $\varphi$, a similar, more careful approach proves much more. We delay the proof of the Section 5.4 below.
Proposition 15
Let $L$ be a displaceable Lagrangian of $M$, and take a compact $K$ containing $L$ in its interior. Suppose furthermore that one of the following holds:
  1. (a)
    $M$ is rational, i.e. $\omega(\pi_{2}(M))$ is discrete, or
  2. (b)
    the image of $\pi_{1}(L)\to\pi_{1}(M)$ is torsion.
Then, there exists $\delta>0$ with the following property. If $\varphi$ is a Hamiltonian diffeomorphism supported in $K$ and $d_{C^{0}}(\mathds{1},\varphi)<\delta$, then $L\cap\varphi(L)\neq\varnothing$.
Furthermore, we get a rigidity result for sequences of Hamiltonian or symplectic diffeomorphisms from Theorem 7 and Corollary 9.
@@
Local contractibility of $\mathcal{L}{\mathrm{Ham}}(L)$ in the Hausdorff topology
While we need the NLC in all cases where we can prove Speculation A, the fact that $\mathcal{L}{\mathrm{Ham}}(L)$ is locally path connected in the Hausdorff topology turns out to be easier to prove. More precisely, we get the following.
@@
Suppose that $L$ is $H$-rational and respects the hypotheses of Theorem 4. Then $\mathcal{L}{\mathrm{Ham}}(L)$ is locally contractible in the Hausdorff topology.
Note that this is not quite the type of results we mentioned earlier in the introduction. Indeed, we do not claim that the Hausdorff-continuous path from $L^{\prime}$ to $L$ (as subsets of the symplectic manifold) in a small neighbourhood of $L$ is generated by an actual Hamiltonian isotopy, simply that it consists at all times of elements of $\mathcal{L}{\mathrm{Ham}}(L)$.
@@
Note that it suffices to prove this statement at $L$. Fix a Weinstein neighbourhood $\Psi:D^{*}_{r}L\to\mathcal{W}(L)$ as given by the conclusions of Theorems 4 and 5. Then, every $L^{\prime}\in\mathcal{L}{\mathrm{Ham}}(L)$ which is in $\mathcal{W}(L)$ is exact in that neighbourhood. We can thus take
@@
to be the contraction. Indeed, exactness in $\mathcal{W}(L)$ ensures that this is a Hamiltonian isotopy for all $t>0.$ In particular $\Psi(t\Psi^{-1}(L^{\prime}))$ is Hamiltonian isotopic to $L^{\prime}$ and hence it is in $\mathcal{L}{\mathrm{Ham}}(L)$. But exactness also implies that the projection $\Psi^{-1}(L^{\prime})\to T^{*}L\to L$ is a homotopy equivalence [AK18]. In particular, that projection must be surjective, otherwise $H_{n}(L^{\prime})\to H_{n}(L)\neq 0$ would be zero. Therefore, $L^{\prime}$ being close to $L$ implies that $L$ is close to $L^{\prime}$ and hence that $L^{\prime}$ and $L$ are Hausdorff-close. This means that the Hausdorff limit of $\Psi(t\Psi^{-1}(L^{\prime}))$ as $t\to 0$ is precisely $L$, i.e. the above contraction is indeed Hausdorff-continuous.
@@
From Theorems 4 and 5, we get the following.
@@
Let $L$ be a $H$-rational Lagrangian in a symplectic manifold, and denote by $\tau$ its $H$-rationality constant. Then, $\mathcal{L}{\mathrm{Symp}_{0}}(L)$ is open in $\mathcal{L}(L,\tau)$ in the $C^{1}$ topology. If moreover $L$ respects the hypotheses of Theorem 4 or $H_{1}(L;\mathds{R})=0$ and the NLC holds on $T^{*}L$, then the same holds in the Hausdorff topology.